OpenStudy (anonymous):

Can anyone help? Write 4 square root a^3b in exponential form?

jimthompson5910 (jim_thompson5910):

this? \[\Large \sqrt[4]{a^3b}\]

OpenStudy (anonymous):

Yes!

jimthompson5910 (jim_thompson5910):

Rule: \[\Large \sqrt[n]{x^m} = (x^m)^{1/n}\]

jimthompson5910 (jim_thompson5910):

so that allows us to simply say \[\Large \sqrt[4]{a^3b} = (a^3b)^{1/4}\] we could go further and rearrange terms but it's in exponential form which is probably what your book wants

jimthompson5910 (jim_thompson5910):

I guess you could add this on \[\Large \sqrt[4]{a^3b} = (a^3b)^{1/4}= a^{3/4}b^{1/4}\] (notice I multiplied the outer exponent by each inner exponent) but it's not really necessary

OpenStudy (anonymous):

Thank you so much for the help. Could you help me with one more?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Simplify 3 square root 2 to the second power

jimthompson5910 (jim_thompson5910):

so this? \[\Large \sqrt[3]{2^2}\]

OpenStudy (anonymous):

Yes, and it's in parenthesis

jimthompson5910 (jim_thompson5910):

try out the rule I gave you and tell me what you get

jimthompson5910 (jim_thompson5910):

alternatively, you can compute 2^2 to get 4, then use the rule

jimthompson5910 (jim_thompson5910):

tell me what you get